Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O4c)
menzione - 22:54 Saturday 13 September 2025 (67694) Print this report
Operator Report - Afternoon shift

ITF found locked at LN3 in SCIENCE mode.
At 19:30 UTC I set CALIBRATION mode and started the planned calibrations:
- 19:30 UTC - CHECK HREC 
- 19:50 UTC - ISC injections

At 20:06 UTC ITF back in SCIENCE mode.

Guard tour (UTC):
13:15 - 14:00
16:24 - 16:59
19:10 - 19:50

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×